Hello,

I need help with this problem:

If 500 sheets of paper weigh 7 lbs, how much do 2 sheets weigh?

Thanks

Answer by checkley75(3666)

7/500=X/2
500X=14
X=14/500
X=.028 LBS FOR THE 2 SHEETS.